A square that has 18 cm long edges is centered on the x axis in a region where there exists a uniform electric field given by E = (1.00 kN/C)i. (a) What is the electric flux of this electric field through the surface of a square if the normal to the surface is in the +x direction? N m2/C (b) What is the electric flux through the same square surface if the normal to the surface makes a 50° angle with the y axis and an angle of 90° with the z axis? N m2/C
Answer #1

here,

sidelength of square , a = 18 cm = 0.18 m

E = 1 KN/C i = 1000 N/C i

a)

the electric flux of this electric feild through the surface , phi = E * a^2 N.m^2 /C

phi = 1000 * 0.18^2 N.m^2/C

phi = 32.4 N.m^2/C

b)

when theta = 50 degree

the electric flux of this electric feild through the surface , phi = E * a^2 * sin(50) N.m^2 /C

phi = 1000 * 0.18^2 * sin(50) N.m^2/C

phi = 24.8 N.m^2/C